### v1

@@ 3.4. The DW-model with manipulative agents
Using the same notation as in previous sections, we can define the iteration for the DW–model with manipulators by
@@
In figures 14 and 15, we show a simulation for the DW–model with the same parameters as in simulation of figures 9 and 10 except the number of iterations until manipulators reach their final opinion, which are chosen $t_{\Delta}\in\{120k\}_{k=1}^{6}$. Observe that we obtain a similar dynamics but with a greater fragmentation in final clusters.
@@
Figure 14. Simultaneous DW–model: Trajectories of an equispaced initial opinion distribution of 100 normal agents and a group of 8 manipulators with a confidence threshold $\varepsilon=0.1$. The manipulative group changes its opinion lineally from $f(0)=-0.8$ to $f(t_{\Delta})=0.8$ for $t_{\Delta}\in\{120k\}_{k=1}^{6}$ iterations. The dashed line represents the trajectory of the opinion of the manipulative group.
Figure 15. Simultaneous DW–model: Clusters with their number of agents. The black arrow and dashed lines indicate the range of the opinion of the manipulative agents corresponding to the dynamics of figure 14.
@@
Next, we will study the DW–model in a similar way as we have done in subsection 3.3.1, but in this case, due to its random component, we will perform 100 simulations for each pair of fixed values $(t_{\Delta},K)$, calculating the corresponding average of the statistical values. It should be noted that medians and median absolute deviations behave, on average, very similarly to means and standard deviations, so we will only show the latter.
The initial distribution of normal agents for all simulations is $\mathbf{x}(0)=\{-1+\frac{2i}{101}\}_{i=1}^{100}$ and, again, we consider a group of $K$ manipulators whose opinion changes linearly from $f(0)=-0.8$ to $f(t_{\Delta})=0.8$ remaining constant $f(t)=0.8$ for $t\geq t_{\Delta}$. The range of $K$ will be the same as for the HK-model: $K\in\{0,\ldots,30\}$ but the values of $t_{\Delta}$ are chosen higher since the convergence is slower. So, we will take $t_{\Delta}\in\{30+30j\}_{j=0}^{30}=\{30,60,90,\ldots,930\}$ and set the stopping criterion of iterations by making it so that, after each iteration, the values of the opinions, rounded to 3 decimal places, are clustered at a distance among clusters greater than the confidence threshold $\varepsilon=0.1$.
In figure 16, it is observed (left) that the mean values increase as the number of the iterations $t_{\Delta}$ (slow change in the opinion of manipulators) and/or the number $K$ of manipulative agents in the group increase. The brightest yellow indicates mean values close to $x=0.72$, which do not reach the value of the final opinion of the manipulative group. This is because there is a certain dispersion (center) in the final distribution as small groups appear outside the range of opinion of the manipulative group. Due to the random nature of the model, the manipulative group is not able to influence, at each iteration, all normal agents in the confidence interval. Note that the mean value of the number of clusters (right) is in the range $[2,11]$, so there is generally no consensus. In fact, in this third graph, the darkest blue represents mean values close to $c=2.5$ clusters. Notice that, in figures 14 and 15, we can observe that, for $t_{\Delta}=720$, the simulation stops with three final clusters but, due to the random character of this model, it is possible to get simulations with a lesser number of final clusters.
@@
Figure 16. Simultaneous DW–model: Means of mean values (left), standard deviation (centre) and number of final clusters (right) after 100 simulations for every number $K$ of manipulators in the group and every number $t_{\Delta}$ of iterations, with a constant confidence threshold $\varepsilon=0.1$.
Regarding the range of $\alpha$of opinions around the median, we can see that when we take $\alpha=90$%, there is an asymmetry in the interval since the subinterval $[x_{6},\mathbf{x}_{M}]$ with values lower than the median is longer than the subinterval $[\mathbf{x}_{M},x_{95}]$ with higher values (see figure 17). But if we consider $\alpha=80$% or lesser, we do not take into account the radical agents not influenced by the manipulative group with opinions close to $x=\pm 1$, and the symmetry appears even when large number of final clusters are obtained.
Figure 17. Simultaneous DW–model: For $\alpha=90$% we represent the length of the subintervals $[x_{6},\mathbf{x}_{M}]$ (left) and $[\mathbf{x}_{M,}x_{95}]$ (centre) and difference between both length (right).
As in the case of the HK-model, in the simultaneous DW-model, the manipulative group moves the mean values of the opinion of the normal agents in the same direction as the manipulators, but whose influence is not so great as to reach consensus on the opinion final value of the manipulators. In fact, the final opinions of normal agents are grouped into a larger number of clusters.
@@ 3.5. The Weighted HK models with manipulative agents
In this section, we will consider the simultaneous Weighted HK models interacting with one group of manipulators. The iterations in these models are analogous to expressions (2.6), (2.7), (2.8), (2.9) and (2.10) except we have to substitute $x_{j}$ by $z_{j}$ where the latter will be an agent (manipulator or normal) in the confidence interval of the normal agent $x_{i}$.
3.5.1. Group and pairwise interactions models
Both the model for group and for pair interactions have a strong tendency to converge, bi-modally, to the opinions in the extremes of the range. This means that a big group of manipulators (even twice the size of the entire normal population) does not alter, at least significantly, the final distribution of opinions. In these models, convergence occurs with a small number of iterations, so if the opinion of the manipulative group opinions varies slowly, the normal agents have enough time to approach the extreme opinions so that the manipulators do not find normal agents in their confidence interval. Only when the number of manipulator is large (the same size as the normal population or larger), they are able to carry some of the few moderate opinions of normal agents to the extremal opinion nearest to the final opinion of manipulators.
3.5.2. Attractive and Repulsive Weighted HK models
@@
In the case of the Simultaneous Attractive Weighted HK model, the opinion dynamics with a manipulative group is richer than for the previous models depending on the values of the threshold $\varepsilon$, the number of iterations $t_{\Delta}$ and the ratio between the number of manipulators in the group and the number of normal agents.
@@
As described in section 2.3, whenever the value of $\varepsilon$ is small (see the graphic at the left of figure 7), the final distribution comprises three major clusters at $x=-1,0,+1$ with some few isolated opinions in between. So, the influence of the manipulative group provokes perturbations on the major clusters and can modify the isolated opinions even grouping them to some cluster. This behaviour is illustrated in figure 18, where a simulation is done with 100 normal opinions initially equispaced and a manipulative group of 60 agents for a confidence threshold $\varepsilon=0.1$. The number of iterations $t_{\Delta}$ to move the opinion of the manipulators from $f(0)=-0.8$ to $f(t_{\Delta})=0.8$ is chosen as $t_{\Delta}\in\{60,120,180\}$.
@@
Figure 18. Trajectories of 100 normal opinions initially equispaced normal agents and a group of 60 manipulators according to the Simultaneous Attractive Weighted HK model. The confidence threshold is fixed to $\varepsilon=0.1$ and the iterations $t_{\Delta}\in\{60,120,180\}$. The initial opinion of manipulators is $f(0)=-0.8$ and the final opinion is $f(t_{\Delta})=0.8$. The trajectory of the opinion of manipulative group follows the dashed line.
@@
For larger values of $\varepsilon$, the distribution converges into two clusters at $x=\pm 1$, with similar number of agents each. Adding a manipulative group, it can move the opinions of normal agents from one side of the opinion space to the other, depending on the size of the group and the trajectory of its opinion. The simulation of figure 19 is made on an equispaced initial distribution of 100 normal agents and a manipulative group of 60 agents for a confidence threshold $\varepsilon=0.6$. Observe that the opinion of manipulators, again, runs from $f(0)=-0.8$ to $f(t_{\Delta})=0.8$ in $t_{\Delta}\in\{10,20,30\}$ iterations. For the value $t_{\Delta}=10$ (left), the final distributions consist of a cluster composed of approximately half of the opinions at $x=-1$ and the rest oscillates between $x=1$ and $x=0.8$ because the manipulators become stubborn within the confidence interval of normal agents close to $x=1$. In the case of $t_{\Delta}=20$ (center), the manipulators move the opinions formerly in the cluster $x=-1$ to a new cluster at $x=0$, and the rest of opinions still oscillate between $x=1$ and $x=0.8$. Finally, for $t_{\Delta}=30$ (right), the manipulators move all the opinions to an only cluster $x=1$. In this case, the number of manipulators is lesser than the number of normal agents (the entire normal population), so the manipulative group is not strong enough to cause the opinion of normal agents to oscillate for many iterations between $x=1$ and $x=0.8$.
@@
Figure 19. Trajectories of an equispaced initial distribution of 100 normal agents and a group of 60 manipulators according to the Simultaneous Attractive Weighted HK model. The confidence threshold is fixed to $\varepsilon=0.6$ and the iterations $t_{\Delta}\in\{10,20,30\}$. The initial opinion of manipulators is $f(0)=-0.8$ and the final opinion is $f(t_{\Delta})=0.8$. The trajectory of the opinion of manipulative group follows the dashed line.
@@
As in section 3.3.1, we will analyse how a manipulative group with $K$ agents affects the opinion dynamics of the Simultaneous Attractive Weighted HK model. We will proceed by fixing the confidence threshold $\varepsilon$ but varying the number of iterations $t_{\Delta}\in\{5j\}_{j=1}^{30}$ and the number of manipulators in the group as $K\in\{5j\}_{j=0}^{30}$. For every pair $(t_{\Delta},K)$, we will do 100 simulations and compute the average values obtained from the mean opinions of normal agents, from the standard deviations and from the number of final opinions (maybe clusters and oscillating or isolated opinions). We have to take into account that, as seen in figures 18 and 19, we can find cases in which the opinion dynamics does not converge and there exists normal agents with oscillating opinions. So, for each simulation, we will consider the opinion distribution in two particular iterations: $t=t_{\Delta}$ where the manipulative group reaches its final opinion and $t=500$ which allows the stabilization of the distribution in case it occurs. For the values $t\in[t_{\Delta},500]$ the manipulative group becomes stubborn.
Figure 20 shows the results for fixed $\varepsilon=0.6$. The upper row corresponds with the values computed at $t=t_{\Delta}$ while the lower row shows the results at $t=500$. The comparative between these two classes of results permits to distinguish the influence of the manipulators to move the opinion of normal agents and their ability (as stubborn group) to retain the opinion of the latter. Observe that the scales for $t_{\Delta}\in[5,150]$ and $K\in[0,150]$ have been chosen larger than for the DW and HK models (cf. sections 3.3 and 3.4) due to the little influence that a small or rapidly changing manipulative group has on the normal population.
In the case of $t=t_{\Delta}$ (upper row) we can observe the following. For $K<30$ with $t_{\Delta}$ greater than $K$, the average of mean opinions is close to $x=0$ (dark blue in the left picture) with a average of standard deviations close to $\sigma=1$ (yellow in the centre) and a small number of final opinions (dark blue in the right). This implies that the opinion distribution is near a bimodal distribution $x=\pm 1$. Whenever $t_{\Delta}<20$, the average of mean opinions is also close to $x=0$ but with smaller dispersion $\sigma\simeq 0.6$ (green in the centre) and a large number of different opinions (yellow at the right picture). From these regions, through intermediate states of transition, one arrives at $t_{\Delta}\gtrsim 40$ and $K\gtrsim 40$, where the manipulators bring the opinion of the normal agents close to their own. Approximately if $K\in[50,100]$ with $K<t_{\Delta}$ (yellow area in the left), the average opinions of normal agents are close to $x=1$ since the number of manipulative agents is not large enough to retain them in their final opinion $x=0.8$. If $K>50$ and $K>t_{\Delta}$ (orange in the left picture), the manipulative group achieves that the average opinion of normal agents is around $x=0.8$, although with some dispersion $\sigma\simeq 0.3$ due to oscillating opinions.
Now, if we compare with the results after $t=500$ iterations (lower row), we can observe that the region with the average of mean opinions $x=1$ (yellow in the left picture) is clearly delimited below $K=100$. This implies that the manipulative group is able to attract normal agents to its final opinion $x=0.8$, but it cannot retain them in that same opinion. It is the fixed point $x=1$ of the model that ends up attracting them. Only when the size of the manipulative group is greater than the total population of normal agents, that is $K>100$, the opinion of the normal agents oscillates around $x=0.8$ obtaining that average value but with dispersion.
@@
Figure 20. Simultaneous Attractive Weighted HK model. After 100 simulations for every numbers $K\in\{5j\}_{j=0}^{30}$ of manipulative agents in the group and $t_{\Delta}\in\{5j\}_{j=1}^{30}$ of iterations that the manipulators are using to vary their initial opinion $f(0)=-0.8$ to their final opinion $f(t_{\Delta})=0.8$, the pictures show the mean values of the mean opinions (left), of the standard deviation (centre) and of number of final opinions (right). The confidence threshold is fixed at $\varepsilon=0.6$ constant.
It is worth mentioning that for small confidence intervals, for example $\varepsilon=0.1$, the influence of the manipulative group at $t=t_{\Delta}$ has an opposite result to what is intended. Notice in the figure 21 that far from moving the opinions of the normal agents towards $x=0.8$, what is achieved is a slight shift in the opposite direction. The blue area in the picture on the left indicates negative values. Also observe that, for all values of $K$ and $t_{\Delta}$ in the simulation, we obtain a moderated dispersion and a moderate or high number of final opinions for $t=t_{\Delta}$.
Figure 21. Simultaneous Attractive Weighted HK model. Simulation with the same parameter as the one in figure 20 except that the confidence threshold is fixed at $\varepsilon=0.1$.
In the case of the Simultaneous Repulsive HK model, recall that it essentially has two behaviours. When the confidence threshold $\varepsilon$ is small, opinions oscillate in the central part of the opinion space so that, after some iterations, some opinions leave that central zone to converge at one of the extremes. If the confidence threshold increases, so does the number of opinions coming out the central zone, so that if $\varepsilon$ is large, the oscillations disappear and normal opinions converge to one of the extremes.

### v2

@@ 4.1. The DW-model with manipulative agents
Using the notation introduced in (2.4), we can define the iteration for the DW–model with manipulators by
@@
The simulation of figure 11 corresponds to the DW–model with 100 equispaced normal agents, one group of $K=8$ manipulators changing theirs opinions from $f(0)=-1$ to $f(t_{\Delta})=1$ for $t_{\Delta}\in\{300,600,900\}$ and $\epsilon=0.1$. Note that we obtain similar dynamics to the HK–model but with more fragmentation in the final clusters.
@@
Figure 11. Simultaneous DW–model: trajectories of an equispaced initial opinion distribution of 100 normal agents and a group of 8 manipulators with a confidence threshold $\varepsilon=0.1$. The manipulative group changes its opinion linearly from $f(0)=-1$ to $f(t_{\Delta})=1$ for $t_{\Delta}\in\{300k\}_{k=1}^{3}$ iterations. The dashed line represents the trajectory of the opinion of the manipulative group.
4.1.1. Simulations for DW–model with one manipulative group
@@
Next, we will study the DW–model in a similar way as we have done in subsection 3.1.2, but in this case, due to its random component, we will perform 100 simulations for each pair of fixed values $(t_{\Delta},K)$, calculating the corresponding central points, dispersion and number of clusters.
@@
Figure 12. Simultaneous DW–model: averages of mean values (left), standard deviations (center) and numbers of final clusters (right) after 100 simulations for every number $K$ of manipulators in the group and every number $t_{\Delta}$ of iterations, with a constant confidence threshold $\varepsilon=0.1$. The opinion of manipulative group varies linearly from $f(0)=-1$ to $f(t_{\Delta})=1$ before to be stubborn $f(t)=1$ for $t\geq t_{\Delta}$.
The initial distribution of normal agents for all simulations is $\mathbf{x}(0)=\{-1+\frac{2i}{101}\}_{i=1}^{100}$ and, again, we consider a group of $K$ manipulators whose opinion changes linearly from $f(0)=-1$ to $f(t_{\Delta})=1$ remaining constant $f(t)=1$ for $t\geq t_{\Delta}$. The range of $K$ will be the same as for the HK-model: $K\in\{0,\ldots,30\}$ but the values of $t_{\Delta}$ are chosen higher since the convergence is slower. So, we will take $t_{\Delta}\in\{30j\}_{j=0}^{30}$ and set the stopping criterion of iterations by making it so that, after each iteration, the values of the opinions, rounded to 3 decimal places, are clustered at a distance among clusters greater than the confidence threshold $\varepsilon=0.1$.
Figure 13. Simultaneous DW–model: averages of centers (left), amplitudes (center) and number of primary clusters (right) after 100 simulations with the same parameters as in figure 12.
In figures 12 and 13, it is observed a very similar result that in the case of HK–model except that the transition region of intermediate values (in greens) in each graph is considerably narrower. This implies that when the manipulative group has significant influence, it manages to attract all normal opinions and achieves the consensus in its own opinion. When $t_{\Delta}\lesssim 300$ or $K\lesssim 8$, the normal population shows some resistance to the influence of manipulators, but when manipulators manage to break that resistance, they get consensus in their opinion. In this case, after observing the effective weights obtained in the simulation (see figure 14), we have used the screening value $\delta=0.2$ to select the primary clusters.
Figure 14. Simultaneous DW–model: histogram of effective weights in simulation of figures 12 and 13.
5. The Weighted HK–models
In order to build an agent–based model to study the influence of fake news on society, Toccaceli et al. [18], Milli [15] define different models as in (2.2) comprising mechanisms of attraction and/or repulsion that we will introduce below. This models include a weight function of the neighbouring opinions [6].
Let us consider an agent $p_{i}$ and a random opinion $x_{j}(t)$ in the distribution chosen such that $j\neq i$. We will call Attractive Weighted HK model to the iteration (5.1) corresponding to the attractive mechanism. The opinion $x_{i}(t)$ is modified by
where $s_{ij}(t)=sign(x_{i}(t)-x_{j}(t))$. The components of the matrix of weights $w=(w_{ij})$ describe the trustworthiness the agent $p_{i}$ attributes to the $p_{j}$’s opinion for $j\in\{j:|x_{i}(t)-x_{j}(t)|\leq\varepsilon\text{ and }j\neq i\}$ and their values will be taken as $w_{ij}\in[0,1]$.
For the model (5.1), if the opinion $x_{j}(t)$ is contained in the confidence interval of $x_{i}(t)$, then there is an attractive influence of $x_{j}(t)$ on the opinion $x_{i}(t)$, that is, the new opinion $x_{i}(t+1)$ moves closer to $x_{j}(t)$.
Note that the term $\phi_{i}(t)=1-|x_{i}(t)|$ can be thought of as the probability density function $g(x)=1-|x|$ that gives weight to the interval $[-1,1]$. This makes moderate opinions (close to $x=0$) more likely to be modified and, therefore, after stabilization of opinion dynamics, most opinions are found in clusters $x=\pm 1$ which are, in fact, fixed points of the iterations. Some opinions may fall outside these radical groups depending on the values of the confidence threshold $\varepsilon$. If the value of $\varepsilon$ is large enough, the final distribution is bimodal, polarized at $x=\pm 1$.
Figure 15. Simultaneous Attractive, Repulsive and Combined Weighted HK models: opinion dynamics for an equispaced initial opinion distribution of $N=100$ normal agents, a threshold of confidence of $\varepsilon=0.1$ and the same random matrix $w=(w_{ij})$.
In the figure 15 and in order to compare the different dynamics, simulations of 100 iterations for the Simultaneous Attractive, Repulsive and combined (Attractive-Repulsive) models are shown for the same equispaced initial opinion distribution of $N=100$ normal agents with a confidence threshold of $\varepsilon=0.1$ and a random weight matrix $w=(w_{ij})$. When the value of $\varepsilon$ is small, the final state for the Attractive model consists of three big clusters and some individual opinions in between. The largest cluster has the opinion of about 45 normal agents and it is located at $x=0$. There are two smaller clusters located at $x=\pm 1$ with about 24 normal agents each one. The opinions of the rest of normal agents are isolated in between those three clusters. For Repulsive and Attractive–Repulsive models, the central opinions oscillate with no convergence.
@@ 5.1. The Weighted HK models with manipulative agents
In this section, we will consider the simultaneous Weighted HK models interacting with one group of manipulators. The iterations in these models are analogous to expressions (5.1), (5.2) and (5.3) except we have to substitute $x_{j}$ by $z_{j}$ where the latter will be an agent (manipulator or normal) in the confidence interval of the normal agent $x_{i}$.
5.1.1. Attractive Weighted HK model with manipulators
@@
The opinion dynamics under Weighted HK models with a manipulative group is richer than for the previous models depending on the values of the threshold $\varepsilon$, the number of iterations $t_{\Delta}$ and the ratio between the number of manipulators in the group and the number of normal agents.
@@
As described in section 5 above, whenever the value of $\varepsilon$ is small (see the graphic at the left of figure 15), the final distribution comprises three major clusters at $x=-1,0,+1$ with some few isolated opinions in between. So, the influence of the manipulative group provokes perturbations on the major clusters and can modify the isolated opinions even grouping them to some cluster but it can not change the opinion of a cluster. This behaviour is illustrated in figure 16, where a simulation is done with 100 normal opinions initially equispaced and a manipulative group of 60 agents for a confidence threshold $\varepsilon=0.1$. The number of iterations $t_{\Delta}$ to move the opinion of the manipulators from $f(0)=-0.9$ to $f(t_{\Delta})=0.9$ is chosen as $t_{\Delta}\in\{60,120,180\}$.
@@
Figure 16. Simultaneous Attractive Weighted HK model: trajectories of 100 normal opinions initially equispaced normal agents and a group of 60 manipulators. The confidence threshold is fixed to $\varepsilon=0.1$ and the iterations $t_{\Delta}\in\{60,120,180\}$. The initial opinion of manipulators is $f(0)=-0.9$ and the final opinion is $f(t_{\Delta})=0.9$. The trajectory of the opinion of manipulative group follows the dashed line.
@@
For larger values of $\varepsilon$, a manipulative group can influence the opinion of clustered normal agents. The simulation of figure 17 is made on the same parameters as in figure 16 except for a confidence threshold $\varepsilon=0.6$ and $t_{\Delta}\in\{10,20,30\}$. Notice that the final clustered opinions change in each graph: $x=\pm 1$ (left), $x=0,1$ (center) and $x=1$ (right). There are oscillating opinions between $x=0.9$ and $x=1$ because the manipulators become stubborn within the confidence interval of normal agents close to $x=1$. For $t_{\Delta}=30$ (right), the manipulators move all the normal opinions to an only cluster $x=1$ and since the number of manipulators is lesser than the number of neighbouring normal agents, the manipulative group is not strong enough to cause the opinion of normal agents to oscillate for many iterations between $x=0.9$ and $x=1$.
@@
Figure 17. Simultaneous Attractive Weighted HK model: trajectories of 100 normal opinions initially equispaced normal agents and a group of 60 manipulators. The confidence threshold is fixed to $\varepsilon=0.6$ and the iterations $t_{\Delta}\in\{10,20,30\}$. The initial opinion of manipulators is $f(0)=-0.9$ and the final opinion is $f(t_{\Delta})=0.9$. The trajectory of the opinion of manipulative group follows the dashed line.
5.1.2. Simulations for Attractive Weighted HK model with one manipulative group
As in section 3.1.2, we will analyse how a manipulative group with $K$ agents affects the opinion dynamics of the simultaneous Attractive Weighted HK model. We will proceed by fixing the confidence threshold to $\varepsilon=0.6$ because the influence of the manipulative group for small values of $\epsilon$ is not significant.
@@ 5.1.2. Simulations for Attractive Weighted HK model with one manipulative group
We will consider an initial equispaced distribution $\mathbf{x}(0)=\{-1+\frac{2i}{101}\}_{i=1}^{100}$ of normal agents and one manipulative group with $K\in\{5j\}_{j=0}^{40}$ agents which change their opinion from $f(0)=-0.9$ to $f(t_{\Delta})=0.9$ for $t_{\Delta}\in\{5j\}_{j=0}^{40}$ and then, for the values $t\in[t_{\Delta},500]$ the manipulative group becomes stubborn. For every pair $(t_{\Delta},K)$, we will do 100 simulations and compute the average values obtained for the central point and width in the two used metrics. We have to take into account that, as seen in figures 16 and 17, we can find cases in which the opinion dynamics does not converge and there exists normal agents with oscillating opinions. So, for each simulation, we will consider the opinion distribution in two particular iterations: $t=t_{\Delta}$ where the manipulative group reaches its final opinion and $t=500$ which allows the stabilization of the distribution in case it occurs. The comparative between these two classes of results permits to distinguish the influence of the manipulators to move the opinion of normal agents and their ability (as stubborn group) to retain the opinion of the latter.
@@
Figure 18. Simultaneous Attractive Weighted HK model: averages of mean opinions (left), of standard deviations (center) and of number of final clusters (right) after 100 simulations for every $K\in\{5j\}_{j=0}^{40}$ and $t_{\Delta}\in\{5j\}_{j=1}^{40}$. The manipulative group varies its opinion from $f(0)=-0.9$ to $f(t_{\Delta})=0.9$. The confidence threshold is fixed at $\varepsilon=0.6$ constant.
Notice that, in figures 18 and 19, the upper row corresponds with the values computed at $t=t_{\Delta}$ while the lower row shows the results at $t=500$. Observe that the scale for $K\in[0,200]$ have been chosen larger than for the DW and HK–models (cf. sections 3.1 and 4.1) due to the little influence that a small manipulative group has on the normal population.
Taking into account the mean and the standard deviation, the influence of the manipulative group is small when $K\lesssim 50$ or $t_{\Delta}\lesssim 25$, since the averages of mean opinions are close to $x=0$ (dark blue in the left pictures). Their influence is important when $K\gtrsim 50$ and $t_{\Delta}\gtrsim 25$ with a narrow transition zone. For this values, the manipulators bring the opinion of the normal agents close to their own. Approximately if $K\in[50,100]$ with $t_{\Delta}\in[100,200]$ (bright yellow area in the left), the average opinions of normal agents are close to $x=1$ since the number of manipulative agents is not large enough to retain them in their final opinion $x=0.9$. If $K>100$ and $t_{\Delta}\in[100,200]$ (dark yellow in the left pictures), the manipulative group achieves that the average opinion of normal agents is around $x=0.9$, although with some dispersion $\sigma\simeq 0.2$ due to oscillating opinions. Moreover, when $K$ is greater than the normal population $N=100$, fragmentation of the opinion occurs.
Let us compare with the results according to the interval of primary clusters (in figure 19). We observe that no more than 3 primary clusters reproduce the same behaviour shown for means and standard deviations. To select the primary clusters we have chosen the filtering value $\delta=0.5$, thus eliminating the oscillating opinions that do not end up clustered. The histograms of effective weights are shown in figure 20.
Figure 19. Simultaneous Attractive Weighted HK model: simulation with the same parameters as in figure 18. The pictures show the average of centers (left), of the amplitudes (center) and of numbers of effective clusters (right).
Figure 20. Simultaneous Attractive Weighted HK model: histograms of effective weights for the simulations of figures 18 and 19 at $t=t_{\Delta}$ and $t=500$.
5.1.3. Repulsive Weighted HK model with manipulators
In the case of the Simultaneous Repulsive HK model, dynamics converge to a bimodal distribution at the extremes $x=\pm 1$ with approximately half of the normal opinions in each value. Only if the confidence threshold $\epsilon$ is small, the moderate opinions can oscillate for some iterations before convergence.